Circular No. 7528

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000er IN PGC 9132 R. Chassagne, Ste. Clotilde, Île de Réunion, reports his discovery of an

apparent supernova (mag 15.1) on CCD images taken with an automated 0.30-m telescope on Nov. 23.83 and 24.78 UT. SN 2000er is located at α = 2^{h}24^{m}32^{s}.54, δ = −58^{o}26′18′′.0 (equinox 2000.0), which is 44′′ east and 16′′ south of the center of PGC 9132. Nothing appears at this position on an image taken by Chassagne on Oct. 7.96 (limiting mag 18) or on the Digital Sky Survey image taken on 1975 Sept. 9 (limiting mag 22).

A. Maury, Observatoire de la Côte d’Azur, writes: “I. Hook, R. Gorki, F. Selman, M. Dennefeld, and I took a spectrum (range 520–900 nm) of SN 2000er with the 3.6-m telescope (+ EFOSC2) at the European Southern Observatory (ESO). The spectrum shows a very blue continuum and the absence of a 615-nm trough; no other feature is visible except a narrow P-Cyg profile at the rest wavelength of He I 587.6-nm. This suggests a possible type-Ib supernova, well before maximum, but more observations with a bluer coverage are needed. The redshift was measured as 0.031.”

A. Clocchiatti, Pontificia Universidad Catolica; and M. Turatto, Osservatorio di Padova, report: “We have observed SN 2000er with the 3.6-m telescope (+ EFOSC2) at ESO on Nov. 26.17 UT. The object looks very peculiar. The preliminary reduction of a spectrum (320–740nm, resolution 1.6 nm) shows a strong continuum that can be fitted redward of 400 nm with a blackbody at 8000 K, but rapidly drops blueward of this wavelength. A large number of narrow features are visible. A narrow unresolved P-Cyg profile has the emission centered at 607 nm and could be either Na I D or He I 587.6-nm. The separation of the components corresponds to an expansion velocity of ∼ 800 km/s. Very broad features are possibly developing. Estimated magnitudes for SN 2000er: V = 16.7, B − V = +0.4, V − R = +0.2; at the distance of the parent galaxy, these correspond to M_{V} = −19.1 (H_{o} = 65 km/Mpc/s).”

CI AQUILAE K. Matsumoto and T. Kato, Kyoto University, report: “R-band CCD

photometry using a 0.25-m telescope reveals that this recurrent nova has apparently entered the final rapid-decline stage. Nightly averaged R_{c} magnitudes: Nov. 19.378 UT, 13.01 ± 0.05; 22.441, 13.68 ± 0.28; 23.401, 15.13 ± 0.61; 25.382, 14.48 ± 0.28. These observations correspond to the orbital phases of 0.60, 0.51, 0.05, and 0.27, respectively, based on our refined eclipse ephemeris. The deep fading on Nov. 23 reflects the primary eclipse.”

2000 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green